# Dirac Equation Solution with Generalized tanh-Shaped Hyperbolic Potential: Application to Charmonium and Bottomonium Mass Spectra

### v1

@@ III.2 Analysis of Charmonium and Bottomonium Mass Spectra
In this part, we mainly focus on the non-relativistic modeling of charmonium and bottomonium systems (See detailed modeling information in S3.A section in Ref. [SM, ]), as it exhibits better agreement with experimental data across both low- and high-energy states. The following analysis delves into specific comparisons and explores the underlying physical implications of these findings.
@@
Figure 1: (a) Non-relativistic calculating mass spectra of charmonium $M_{c\tilde{c}}$ in GeV for $m_{c}$ = 1.27 GeV; $V_{1}$ = 7.404662 GeV; $V_{2}$ = -20.779639 GeV; $V_{3}$ = 14.959636 GeV; $\alpha$ = 0.223612 GeV; $r_{e}$ = 3.830931 GeV-1. The experimental data from Ref. [Particle Data Group et al., 2022] and other theoretical predictions: Theory I from Ref. [Purohit et al., 2022], Theory II from Ref. [Soni et al., 2018], Theory III from Ref. [Bukor and Tekel, 2023] are also included for comparison. Corresponding deviations from the experimental data.
Figure 2: Non-relativistic calculating mass spectra of bottomonium $M_{b\tilde{b}}$ in GeV for $m_{b}=4.18$ GeV; $V_{1}$ = 4.883229 GeV; $V_{2}$ = -12.928267 GeV; $V_{3}$ = 10.108982 GeV; $\alpha$ = 0.412272 GeV; $r_{e}$ = 1.836733 GeV-1. The experimental data from Ref. [Particle Data Group et al., 2022] and other theoretical predictions: Theory I from Ref. [Purohit et al., 2022], Theory II from Ref. [Soni et al., 2018], Theory III from Ref. [Bukor and Tekel, 2023] are also included for comparison. Corresponding deviations from the experimental data.
@@
The calculated charmonium mass spectrum, as depicted in Figure 1 and detailed in Table S1 in Ref. [SM, ] and, shows a well agreement with experimental results, particularly for the $1S$ ($J/\psi(1S)$), $2S$ ($\psi(2S)$), and $3S$ ($\psi(4040)$) states. Notably, the $J/\psi(1S)$ state matches precisely at $3.096900$ GeV, underscoring the efficacy of the GTHP model in predicting ground state masses where non-relativistic approximations hold strong.
@@
However, the model’s performance varies across different orbital states. While $S$-wave states are well-described, higher orbital states, such as the $1P$ ($\chi_{c1}(1P)$) and $1D$ ($\psi(3770)$) states, exhibit deviations of -8.96% and -10.51%, respectively. Such discrepancies indicate the need for further refinement in modeling higher orbital angular momentum states. The significant deviations observed in these states suggest that the influence of relativistic effects, spin-orbit coupling, and other acceptable structure corrections become more pronounced. Again, in our context of relativistic treatment, using the Eq. (S32), we observe relatively larger deviations from experimental data, especially in the $1P$, $1D$, and higher orbital states, primarily due to the overestimation of relativistic contributions without proper spin-orbit interaction corrections.
@@
In order to remove these discrepancies these discrepancies, spin symmetry breaking can be introduced through the parameter $C_{s}$, which allows for fine-tuning to match the ground state with experimental values. Although this improved the overall agreement, particularly for the charmonium system, the deviations for higher angular momentum states persisted. This is due to the fact that the Eq. (S32) are obtained for spin symmetry case. The interaction potential may need to be adjusted to account for quark-antiquark dynamics more comprehensively, particularly in exact spin-orbit correction that naturally arise in systems with non-negligible quark masses. To gain further improvements in the relativistic model, one has to go to higher order corrections like Tensor potentials or solve two body DE and Bethe-Salpeter formalism, as these methods account more accurately for spin-dependent interactions in higher excited states. For a detailed exploration of the relativistic mass spectra and a comprehensive comparison between non-relativistic and relativistic results with the full set of mass spectra tables, analysis of spin symmetry breaking, readers can be found in section S3.B of Ref. [SM, ].
Interestingly, the GTHP model’s ability to accurately predict $S$-wave states suggests that the primary contributions to the binding energies in these states are well-captured by the non-relativistic potential. Furthermore, the GTHP model’s framework implies that certain higher excited states, particularly those with significant orbital angular momentum, may exist in more tightly bound configurations than predicted by models incorporating only linear or Coulombic potentials.Purohit et al. (2022); Barnes et al. (2005) This could lead to the discovery of new states with unexpectedly high binding energies, which remain stable under conditions where conventional quarkonium states would dissociate.
@@
For bottomonium, the mass spectra shown in Figure 2 and Table S2 in Ref. [SM, ] reveal an even closer agreement with experimental data, particularly for the $1S$ ($\Upsilon(1S)$), $2S$ ($\Upsilon(2S)$), and $3S$ ($\Upsilon(3S)$) states, where deviations from experimental values are around 0.0%. This demonstrates the robustness of the GTHP model in predicting bottomonium states, where the non-relativistic approximation is more justified due to the larger bottom quark mass. For higher states, such as the $1P$ ($h_{b}(1P)$), $1D$ ($\Upsilon_{2}(1D)$), $2P$ ($h_{b}(2P)$), and $4S$ ($\Upsilon(4S)$) states, the deviations are slightly more significant, ranging from -1.06% to 3.31%. These minor discrepancies indicate that while the GTHP model is highly reliable even for excited states, there is room for improvement by incorporating relativistic and spin-orbital corrections. The model’s accuracy in these states implies that the essential physics of quark confinement and interaction potentials are well captured. However, spin-spin, spin-orbit, and tensor interactions might need finer adjustments for higher precision.
@@
While the GTHP model demonstrates good performance across a range of quantum states, a comparison with other theoretical models as Theory I (non-relativistic analyses of the linear plus modified Yukawa potential)Purohit et al. (2022), Theory II (non-relativistic analyses of quark-antiquark Cornell potential) Soni et al. (2018), and Theory III (non-relativistic analyses of Cornell potential) Bukor and Tekel (2023)-reveals varying degrees of agreement. Although Theory II and III show closer results for lower energy states, the GTHP model provides better accuracy overall, especially for higher orbital states. This indicates that the GTHP has provided a more powerful description of quarkonium systems, especially in the high-energy region. Also, our non-relativistic quantum-mechanical treatment yields important insights, especially in the lower-energy region, but we have to be aware of its limitations when considering higher orbital states or systems with significant relativistic effects. This framework includes spin effects semi-inclusively, but a completely relativistic treatment will necessitate in addition the formulation and solution of two-body DE Crater et al. (2009); Scott et al. (1992); Ferreira (1988) as well as the Bethe-Salpeter equation Hilger et al. (2015). Implicitly in this approach would be the relativistic corrections our model now "corrects", providing a more realistic picture of quark-antiquark dynamics, especially regarding higher angular-momentum states. Note that a relativistic treatment could entail more than just this pair term – in particular, we might have important correlations through spin-orbit coupling, which are known to induce fine structure splittings not accounted for by our model.Particle Data Group et al. (2022)
@@
Such the two-body DE Crater et al. (2009); Scott et al. (1992); Ferreira (1988) and the Bethe-Salpeter equation Hilger et al. (2015) formalism can also open avenues with GTHP for exploring new quarkonium states that may emerge owing to relativistic effects, which are not predicted by conventional non-relativistic models. These heavily bound states may include, e.g., the very tightly-bound ones with large binding energies where tests of the model can be made against experiments; and new states which are created under extremely severe conditions in such things as a heavy-ion collision plasma situations where non-relativistic models might be insufficient. The predictions by this model of possible new states with longevity at extreme conditions can also be significant in the context of heavy-ion collisions, where quark-gluon plasma might be created and where new bound states might be discovered. Expanding this framework and applying it to mixed-flavor systems like bottom-charm ($B_{c}$) mesons can also yield deeper insights into quark confinement and the dynamics of the strong force. Such predictions, although speculative, highlight the need for closing elliptical states not only in the GTHP model but also by including more sophisticated interactions to study a whole range of quarkonium states. Given these, while hardy in its current form, the GTHP model may serve as a starting block for more advanced models. A more accurate treatment together with incorporating more complex interactions will enable a deeper study of quarkonium states and give us access to a wider domain of the strong interactions that are responsible for these configurations.
@@ IV Conclusions
In this study, the analytical solution of the Dirac equation for GTHP was carried out by using the NU method in the context of spin and pseudospin symmetries. The resulting energy eigenvalues were expressed for various quantum states by means of hypergeometric polynomials. The sensitivity of energy eigenvalues to potential parameters underlines the adaptability and versatility of the GTHP framework for modelling complex quantum systems. For charmonium and bottomonium mass spectra analyses, our findings in low-lying $S$-wave states, namely $J/\psi(1S)$, $\psi(2S)$, and $\psi(4040)$, showed a good agreement with reported experimental data, particularly for ground states. This finding highlights the model’s accuracy in the non-relativistic regime, especially for heavy quark systems like bottomonium, where deviations from experimental masses remained minimal. However, in higher orbital states, including $1P$ ($\chi_{c1}(1P)$) and $1D$ ($\psi(3770)$) for charmonium and $1P$ ($h_{b}(1P)$) and $1D$ ($\Upsilon_{2}(1D)$) for bottomonium, relativly larger discrepancies observed. These deviations indicate the increasing importance of spin-orbit coupling and relativistic corrections, especially in states with higher angular momentum. While non-relativistic analyses of the GTHP model adequately capture the fundamental dynamics of heavy quarkonia, it has also been shown that including relativistic and spin-dependent effects is essential to advance the description of higher orbital states. A similar framework for systems with a mixture of flavors, like $B_{c}$ mesons, incorporating spin-orbit corrections can be insightful in the case of quark confinement and strong interaction. Overall, the GTHP model provides a solid foundation for theoretical exploration and reliable predictions, presenting substantial potential for future advancements in the field.
@@ Acknowledgments
Notes and references
@@
[…]
[findent=2pt]Within the framework of advanced theoretical physics, the DE provides a profound description of particles possessing mass $M$ under the influence of an attractive scalar potential $S(\vec{{r}})$ and a repulsive vector potential $V(\vec{{r}})$. This formulation is crucial for understanding the quantum behaviors of fermions in various potentials and is given by Greiner (2001):
@@
$k=\begin{cases}-\left(l+1\right)=-\left(j+\frac{1}{2}\right),&\text{for aligned spin states}\ (k<0),\\ +l=+\left(j+\frac{1}{2}\right),&\text{for unaligned spin states}\ (k>0),\end{cases}$ (S4)
@@
$k=\begin{cases}-\tilde{l}=-\left(j+\frac{1}{2}\right),&\text{foraligned spin states}\ (k<0),\\ +\left(\tilde{l}+1\right)=+\left(j+\frac{1}{2}\right),&\text{for unaligned spin states}\ (k>0),\end{cases}$ (S5)
@@
$\begin{array}[]{l}\left[\frac{d^{2}}{dr^{2}}-\frac{k(k+1)}{r^{2}}-\frac{(Mc^{2}+E_{nk}-\Delta(r))(Mc^{2}-E_{nk}+\Sigma(r))}{{\hslash}^{2}c^{2}}+\frac{\frac{d\Delta(r)}{dr}\left(\frac{d}{dr}+\frac{k}{r}\right)}{Mc^{2}+E_{nk}-\Delta(r)}\right]F_{nk}(r)=0,\\ \\ \left[\frac{d^{2}}{dr^{2}}-\frac{k(k-1)}{r^{2}}-\frac{(Mc^{2}+E_{nk}-\Delta(r))(Mc^{2}-E_{nk}+\Sigma(r))}{{\hslash}^{2}c^{2}}-\frac{\frac{d\Sigma(r)}{dr}\left(\frac{d}{dr}-\frac{k}{r}\right)}{Mc^{2}-E_{nk}+\Sigma(r)}\right]G_{nk}(r)=0,\end{array}$ (S9)
@@
The regime of spin symmetry is delineated by the condition $\frac{d\triangle(r)}{dr}=0$, leading to a constant $\triangle(r)=C_{s}$. This pivotal assumption simplifies the analysis significantly, as delineated in the seminal works cited herein.Ginocchio (2005); Liang et al. (2015) Within this framework, the modified potential terms, $S(r)=V(r)-C_{s}$ and $\Sigma(r)=2V(r)-C_{s}$, are instrumental in reformulating the second equation of our foundational Eq. (S8) set as follows:
@@
where $k=l$ for positive $k$ values and $k=-(l+1)$ for negative ones. The energy levels $E_{nk}$ are functions of the quantum numbers $n$ and $l$, intricately linked to the spin symmetry quantum number. It is noteworthy that non-zero $\tilde{l}$ values engender degenerate states characterized by $j=l\pm{1}/{2}$, epitomizing $SU(2)$ spin symmetry.
@@
Given the spin-orbit centrifugal term $V_{soc}(r)=\frac{k(k+1)}{r^{2}}$, analytical solutions to Eq. (S12) are elusive except for the cases where $k=0$ or $k=-1$. Thus, the widely embraced Pekeris approximation Pekeris (1934) is adopted for the pragmatic resolution of the equation, facilitating the treatment of the spin-orbit centrifugal term Ahmadov et al. (2021); Badalov and Badalov (2023).
@@
The coefficients $A_{0}$, $A_{1}$, and $A_{2}$ are determined through the relations presented below, as derived in the Ref.[Ahmadov et al., 2021; Badalov and Badalov, 2023]:
@@
Utilizing these expressions and following the methodology outlined in Ref.[Nikiforov and Uvarov, 1988], we deduce the form of the function $\pi(z)$ by substituting from Eq. (S19) and considering the derivative $\sigma^{\prime}(z)=-2z$, thus obtaining:
@@
Further, by employing the relations for $\lambda=k+\pi^{\prime}(z)$ and the quantized eigenvalue condition $\lambda_{n}=-n\tau^{\prime}(z)-\frac{n(n-1)}{2}\sigma^{\prime\prime}(z)$, where $n=0,1,2,\ldots$, as prescribed in Ref.[Nikiforov and Uvarov, 1988], we acquire:
@@
$\begin{split}&{\left(Mc^{2}-C_{s}\right)}^{2}-E^{2}_{nk}+\left(Mc^{2}+E_{nk}-C_{s}\right)\left(V_{1}+V_{2}+V_{3}\right)\\ &\quad+\frac{{\hslash}^{2}c^{2}k(k+1)}{r^{2}_{e}}\left(A_{0}+A_{1}+A_{2}\right)\\ &\quad={\alpha}^{2}{\hslash}^{2}c^{2}\left(\sqrt{\frac{1}{4}+\frac{\left(Mc^{2}+E_{nk}-C_{s}\right)V_{3}}{{\alpha}^{2}{\hslash}^{2}c^{2}}+\frac{k(k+1)}{{\alpha}^{2}r^{2}_{e}}A_{2}}-n-\frac{1}{2}\right.\\ &\quad\left.+\frac{\frac{\left(Mc^{2}+E_{nk}-C_{s}\right)V_{2}}{2{\alpha}^{2}{\hslash}^{2}c^{2}}+\frac{k(k+1)}{2{\alpha}^{2}r^{2}_{e}}A_{1}}{\sqrt{\frac{1}{4}+\frac{\left(Mc^{2}+E_{nk}-C_{s}\right)V_{3}}{{\alpha}^{2}{\hslash}^{2}c^{2}}+\frac{k(k+1)}{{\alpha}^{2}r^{2}_{e}}A_{2}}-n-\frac{1}{2}}\right)^{2}\end{split}$ (S32)
Here, $n$ takes values from 00 to $n_{max}$ defined as:
@@
The other part of the wave function $y_{n}\left(z\right)$ takes on a hypergeometric-type function whose polynomial solutions are given by the Rodrigues relation Nikiforov and Uvarov (1988). It can be expressed as:
@@
where $C_{nk}$ represents the normalization constant Abramowitz and Stegun (2012).
@@
[…]
@@
Pseudospin symmetry is characterized by the condition $\frac{d\Sigma(r)}{dr}=0$, resulting in $\Sigma(r)=C_{ps}=const$Ginocchio (2005); Liang et al. (2015). This leads to the simplification of the second equation presented in Eq. (S9) set with $S(r)=C_{ps}-V(r)$ and $\triangle(r)=2V(r)-C_{ps}$, yielding:
@@
In cases where the pseudo-centrifugal term ($V_{pc}(r)=\frac{k(k-1)}{r^{2}}$) cannot be solved analytically for arbitrary $k$, the Pekeris approximation scheme Pekeris (1934) Eq. (S13) is applied to it, resulting in the following equation:

### v4

@@ III.2 Analysis of Charmonium and Bottomonium Mass Spectra
In this part, we mainly focus on the non-relativistic modeling of charmonium and bottomonium systems (See detailed modeling information in S3.A section in Ref. [79]), as it exhibits better agreement with experimental data across both low- and high-energy states. The following analysis delves into specific comparisons and explores the underlying physical implications of these findings.
@@
Figure 1: (a) Non-relativistic calculating mass spectra of charmonium $M_{c\tilde{c}}$ in GeV for $m_{c}$ = 1.27 GeV; $V_{1}$ = 7.404662 GeV; $V_{2}$ = -20.779639 GeV; $V_{3}$ = 14.959636 GeV; $\alpha$ = 0.223612 GeV; $r_{e}$ = 3.830931 GeV-1. The experimental data from Ref. [75] and other theoretical predictions: Theory I from Ref. [61], Theory II from Ref. [62], Theory III from Ref. [63] are also included for comparison. (b) Corresponding deviations from the experimental data.
Figure 2: Comparison between the GTHP potential (solid) and a linear confinement potential (dashed). The GTHP realistically captures QCD string-breaking by saturating at large distances, accurately describing compact $S$-wave states but reducing confinement for extended $P$-wave states [85, 86, 87, 88].
@@
The calculated charmonium mass spectrum, as depicted in Figure 1 and detailed in Table S1 and also Table S1 in Ref. [79] and, shows a well agreement with experimental results, particularly for the $1S$ ($J/\psi(1S)$), $2S$ ($\psi(2S)$), and $3S$ ($\psi(4040)$) states. Notably, the $J/\psi(1S)$ state matches precisely at $3.096900$ GeV, underscoring the efficacy of the GTHP model in predicting ground state masses where non-relativistic approximations hold strong.
@@
However, the model’s performance varies across different orbital states. While $S$-wave states are well-described, higher orbital states, such as the $1P$ ($\chi_{c1}(1P)$) state, exhibit deviations of -8.96% (see Figure 1 (a)) for non-relativistic case, respectively. Such discrepancies indicate the need for further refinement in modeling higher orbital angular momentum states. The significant deviations observed in these states suggest that the influence of relativistic effects, spin-orbit coupling, and other acceptable structure corrections become more pronounced. Again, in our context of relativistic treatment, using the Eq. (S32), we observe relatively larger deviations from experimental data, especially in the $1P$ and other higher orbital states, primarily due to the overestimation of relativistic contributions without proper spin-orbit interaction corrections.
Figure 3: (a) Non-relativistic calculating mass spectra of bottomonium $M_{b\tilde{b}}$ in GeV for $m_{b}=4.18$ GeV; $V_{1}$ = 4.883229 GeV; $V_{2}$ = -12.928267 GeV; $V_{3}$ = 10.108982 GeV; $\alpha$ = 0.412272 GeV; $r_{e}$ = 1.836733 GeV-1. The experimental data from Ref. [75] and other theoretical predictions: Theory I from Ref. [61], Theory II from Ref. [62], Theory III from Ref. [63] are also included for comparison. (b) Corresponding deviations from the experimental data.
Another reason for the observed deviations can be that the analytic heavy-$Q\bar{Q}$ potential behaves like a Coulomb+linear (“Cornell”) potential at short distances but flattens at large $r$, thereby mimicking QCD string breaking. Figure 2 compares the GTHP (solid curve) to a conventional linear confining potential. At large separations the GTHP flattens, reflecting the string-breaking screening expected in full QCD [85, 89, 86, 87, 88]. In practice this means that beyond roughly $r\sim 1.0$-$1.2$ fm the effective confining force weakens (approaching the threshold of two non-interacting heavy-light mesons [88]). Consequently, compact $S$-wave quarkonia (with small radii) reside almost entirely within the steep central region of the potential and are well-described by the model. By contrast, $P$-wave states have larger spatial extent and sample the softened tail of the potential; since the confining strength is weaker at these distances, the predicted $P$-wave binding is insufficient. This under-binding of the $P$-waves is therefore a natural consequence of the screened potential choice. Lattice QCD results confirm that the static heavy-quark potential saturates at long range due to light-quark pair creation [86, 88]. We conclude that while the GTHP provides an accurate description of the compact $S$-states, further refinements (e.g. additional interaction terms or coupled-channel effects) will be required to capture the physics of the more extended $P$-wave states.
@@
In order to remove these discrepancies, spin symmetry breaking can be introduced through the parameter $C_{s}$, which allows for fine-tuning to match the ground state with experimental values. Although this improved the overall agreement, particularly for the charmonium system, the deviations for higher angular momentum states persisted. This is due to the fact that the Eq. (S32) are obtained for spin symmetry case. The interaction potential may need to be adjusted to account for quark-antiquark dynamics more comprehensively, particularly in exact spin-orbit correction that naturally arise in systems with non-negligible quark masses. To gain further improvements in the relativistic model, one has to go to higher order corrections like refined GTHP with Tensor potentials or solve two body DE and Bethe-Salpeter formalism with GTHP, as these methods account more accurately for spin-dependent interactions in higher excited states. For a detailed exploration of the relativistic mass spectra and a comprehensive comparison between non-relativistic and relativistic results with the full set of mass spectra tables, analysis of spin symmetry breaking, readers can be found in section S3.B of Ref. [79].
@@
For bottomonium, the mass spectra shown in Figure 3 and detailed in Table S2 and also Table S2 in Ref. [79] reveal an even closer agreement with experimental data, particularly for the $1S$ ($\Upsilon(1S)$), $2S$ ($\Upsilon(2S)$), and $3S$ ($\Upsilon(3S)$) states, where deviations from experimental values are around 0.0%, see Figure . This demonstrates the robustness of the GTHP model in predicting bottomonium states, where the non-relativistic approximation is more justified due to the larger bottom quark mass. For higher states, such as the $1P$ ($h_{b}(1P)$), $1D$ ($\Upsilon_{2}(1D)$), $2P$ ($h_{b}(2P)$), and $4S$ ($\Upsilon(4S)$) states, the deviations are slightly more significant, ranging from -1.06% to 3.31% (see Figure 3(b)). These minor discrepancies indicate that while the GTHP model is highly reliable even for excited states, there is room for improvement by incorporating relativistic and spin-orbital corrections. The model’s accuracy in these states implies that the essential physics of quark confinement and interaction potentials are well captured. However, spin-spin, spin-orbit, and tensor interactions might need finer adjustments for higher precision.
@@
While the GTHP model demonstrates good performance across a range of quantum states, a comparison with other theoretical models as Theory I (non-relativistic analyses of the linear plus modified Yukawa potential) [61], Theory II (non-relativistic analyses of quark-antiquark Cornell potential) [62], and Theory III (non-relativistic analyses of Cornell potential) [63]-reveals varying degrees of agreement. Although Theory II and III show closer results for lower energy states, the GTHP model provides better accuracy overall, especially for higher orbital states. This indicates that the GTHP has provided a more powerful description of quarkonium systems, especially in the high-energy region. Also, our non-relativistic quantum-mechanical treatment yields important insights, especially in the lower-energy region, but we have to be aware of its limitations when considering higher orbital states or systems with significant relativistic effects. This framework includes spin effects semi-inclusively, but a completely relativistic treatment will necessitate in addition the formulation and solution of two-body DE [90, 91, 92] as well as the Bethe-Salpeter equation [93]. Implicitly in this approach would be the relativistic corrections our model now "corrects", providing a more realistic picture of quark-antiquark dynamics, especially regarding higher angular-momentum states. Note that a relativistic treatment could entail more than just this pair term – in particular, we might have important correlations through spin-orbit coupling, which are known to induce fine structure splittings not accounted for by our model [75].
@@
Such the two-body DE [90, 91, 92] and the Bethe-Salpeter equation [66, 93] formalism can also open avenues with GTHP for exploring new quarkonium states that may emerge owing to relativistic effects, which are not predicted by conventional non-relativistic models. These heavily bound states may include, e.g., the very tightly-bound ones with large binding energies where tests of the model can be made against experiments; and new states which are created under extremely severe conditions in such things as a heavy-ion collision plasma situations where non-relativistic models might be insufficient. The predictions by this model of possible new states with longevity at extreme conditions can also be significant in the context of heavy-ion collisions, where quark-gluon plasma might be created and where new bound states might be discovered. Expanding this framework and applying it to mixed-flavor systems like bottom-charm ($B_{c}$) mesons can also yield deeper insights into quark confinement and the dynamics of the strong force. Such predictions, although speculative, highlight the need for closing elliptical states not only in the GTHP model but also by including more sophisticated interactions to study a whole range of quarkonium states. Given these, while hardy in its current form, the GTHP model may serve as a starting block for more advanced models. A more accurate treatment together with incorporating more complex interactions will enable a deeper study of quarkonium states and give us access to a wider domain of the strong interactions that are responsible for these configurations.
@@ IV Conclusions
In this study, we derived the analytical solution of the DE in the context of spin and pseudospin symmetries for the GTHP by using NU method and demonstrated its effectiveness in reproducing the mass spectra of charmonium and bottomonium with sufficient accuracy. The flexibility of the GTHP allows its application in different quantum systems, potentially holding its place in our approach to solving complex fermionic systems. For the analysis of the mass spectra of charmonium and bottomonium, our findings for the low $S$-wave states, namely $J/\psi(1S)$, $\psi(2S)$ and $\psi(4040)$, were in good agreement with the reported experimental data. This finding emphasized the accuracy of the model in the non-relativistic regime, especially for heavy quark systems such as bottomonium, where deviations from experimental masses are minimal. However, in the higher orbital states, such as $1P\longmapsto 4P$ ($\chi_{c1}$) for charmonium and $1P$ ($h_{b}(1P)$) and $1D$ ($\Upsilon_{2}(1D)$) for bottomonium, we observed the relatively larger discrepancy. While the non-relativistic analyses of the GTHP model adequately capture the fundamental dynamics of the heavy quarkonium, it was also shown that the inclusion of relativistic and spin-dependent effects is necessary to advance the description of higher orbital states. For systems with more than one flavour species, such as $B_{c}$ mesons, a corresponding formulation taking spin-orbit corrections into account can shed more light on the propagation of aberrations owing to quark confinement and the strong force. Its overall performance provided a strong basis for theoretical considerations and reliable predictions with significant potential for further development.
@@ Acknowledgments
Data Availability Statement
All data that support the findings of this study are included within the article (and any supplementary files).
Conflict of Interest
The authors declare no competing interests. All research has been carried out within an appropriate ethical framework.
@@
Within the framework of advanced theoretical physics, the DE provides a profound description of particles possessing mass $M$ under the influence of an attractive scalar potential $S(\vec{{r}})$ and a repulsive vector potential $V(\vec{{r}})$. This formulation is crucial for understanding the quantum behaviors of fermions in various potentials and is given by [77]:
@@
[…]
@@
[…]
@@
$\begin{array}[]{l}\left[\frac{d^{2}}{dr^{2}}-\frac{k(k+1)}{r^{2}}-\frac{(Mc^{2}+E_{nk}-\Delta(r))(Mc^{2}-E_{nk}+\Sigma(r))}{{\hslash}^{2}c^{2}}+\frac{\frac{d\Delta(r)}{dr}\left(\frac{d}{dr}+\frac{k}{r}\right)}{Mc^{2}+E_{nk}-\Delta(r)}\right]F_{nk}(r)=0,\\ \\ \left[\frac{d^{2}}{dr^{2}}-\frac{k(k-1)}{r^{2}}-\frac{(Mc^{2}+E_{nk}-\Delta(r))(Mc^{2}-E_{nk}+\Sigma(r))}{{\hslash}^{2}c^{2}}-\frac{\frac{d\Sigma(r)}{dr}\left(\frac{d}{dr}-\frac{k}{r}\right)}{Mc^{2}-E_{nk}+\Sigma(r)}\right]G_{nk}(r)=0.\end{array}$ (S9)
@@
The regime of spin symmetry is delineated by the condition $\frac{d\triangle(r)}{dr}=0$, leading to a constant $\triangle(r)=C_{s}$. This pivotal assumption simplifies the analysis significantly, as delineated in the seminal works cited herein [21, 22]. Within this framework, the modified potential terms, $S(r)=V(r)-C_{s}$ and $\Sigma(r)=2V(r)-C_{s}$, are instrumental in reformulating the second equation of our foundational Eq. (S8) set as follows:
@@
where $k=l$ for positive $k$ values and k=(l+1)k=-(l+1) for negative ones. The energy levels $E_{nk}$ are functions of the quantum numbers $n$ and $l$, intricately linked to the spin symmetry quantum number. It is noteworthy that non-zero $\tilde{l}$ values engender degenerate states characterized by $j=l\pm{1}/{2}$, epitomizing $SU(2)$ spin symmetry.
@@
Given the spin-orbit centrifugal term $V_{soc}(r)=\frac{k(k+1)}{r^{2}}$, analytical solutions to Eq. (S12) are elusive except for the cases where $k=0$ or k=1k=-1. Thus, the widely embraced Pekeris approximation [94] is adopted for the pragmatic resolution of the equation, facilitating the treatment of the spin-orbit centrifugal term [43, 44].
@@
The coefficients $A_{0}$, $A_{1}$, and $A_{2}$ are determined through the relations presented below, as derived in the Ref. [43, 44]:
@@
Utilizing these expressions and following the methodology outlined in Ref. [78], we deduce the form of the function $\pi(z)$ by substituting from Eq. (S19) and considering the derivative $\sigma^{\prime}(z)=-2z$, thus obtaining:
@@
Further, by employing the relations for $\lambda=k+\pi^{\prime}(z)$ and the quantized eigenvalue condition $\lambda_{n}=-n\tau^{\prime}(z)-\frac{n(n-1)}{2}\sigma^{\prime\prime}(z)$, where $n=0,1,2,\ldots$, as prescribed in Ref.[78], we acquire:
@@
$\begin{split}&{\left(Mc^{2}-C_{s}\right)}^{2}-E^{2}_{nk}+\left(Mc^{2}+E_{nk}-C_{s}\right)\left(V_{1}+V_{2}+V_{3}\right)\\ &\quad+\frac{{\hslash}^{2}c^{2}k(k+1)}{r^{2}_{e}}\left(A_{0}+A_{1}+A_{2}\right)\\ &\quad={\alpha}^{2}{\hslash}^{2}c^{2}\left(\sqrt{\frac{1}{4}+\frac{\left(Mc^{2}+E_{nk}-C_{s}\right)V_{3}}{{\alpha}^{2}{\hslash}^{2}c^{2}}+\frac{k(k+1)}{{\alpha}^{2}r^{2}_{e}}A_{2}}-n-\frac{1}{2}\right.\\ &\quad\left.+\frac{\frac{\left(Mc^{2}+E_{nk}-C_{s}\right)V_{2}}{2{\alpha}^{2}{\hslash}^{2}c^{2}}+\frac{k(k+1)}{2{\alpha}^{2}r^{2}_{e}}A_{1}}{\sqrt{\frac{1}{4}+\frac{\left(Mc^{2}+E_{nk}-C_{s}\right)V_{3}}{{\alpha}^{2}{\hslash}^{2}c^{2}}+\frac{k(k+1)}{{\alpha}^{2}r^{2}_{e}}A_{2}}-n-\frac{1}{2}}\right)^{2}.\end{split}$ (S32)
Here, $n$ takes values from 0 to $n_{max}$ defined as:
@@
The other part of the wave function $y_{n}\left(z\right)$ takes on a hypergeometric-type function whose polynomial solutions are given by the Rodrigues relation [78]. It can be expressed as:
@@
where $C_{nk}$ represents the normalization constant [95].
@@
$\int^{\infty}_{0}{{\left|F_{nk}\left(r\right)\right|}^{2}dr}=\frac{1}{\alpha}\int^{1}_{0}{\frac{{\left|F_{nk}\left(z\right)\right|}^{2}dz}{1-z^{2}}}=1\ .$ (S39)
@@
Pseudospin symmetry is characterized by the condition $\frac{d\Sigma(r)}{dr}=0$, resulting in $\Sigma(r)=C_{ps}=const$[21, 22]. This leads to the simplification of the second equation presented in Eq. (S9) set with $S(r)=C_{ps}-V(r)$ and $\triangle(r)=2V(r)-C_{ps}$, yielding:
@@
In cases where the pseudo-centrifugal term ($V_{pc}(r)=\frac{k(k-1)}{r^{2}}$) cannot be solved analytically for arbitrary $k$, the Pekeris approximation scheme [94] Eq. (S13) is applied to it, resulting in the following equation: